OpenStudy (anonymous):

will give medal and fan for best help. I have tried this problem time and time again and i can't get the right answer. please help. (posting equation below)

OpenStudy (anonymous):

\[3d+d-7=\frac{ 25 }{ 4 }\]

OpenStudy (jtug6):

Ok so first things first, what is 3d + d equal to?

OpenStudy (anonymous):

4d

OpenStudy (jtug6):

right so now you got 4d - 7 = 25/4 now add 7 across so you get 4d = 25/4 + 7

OpenStudy (jtug6):

How can you simplify 25/4 + 7?

OpenStudy (anonymous):

it would be \[13\frac{ 1 }{ 4 }\]

OpenStudy (jtug6):

Right so now you have 4d = 13 1/4 or thats the same as 53/4 so ill sticl with 53/4 4d = 53/4 so how do you get d by itsself?

OpenStudy (anonymous):

divide by 4

OpenStudy (jtug6):

yes! so d = 53/4 / 4 which is equal to?

OpenStudy (jtug6):

\[53/4 \div 4\] to make it look prettier :P

OpenStudy (anonymous):

3 5/16

OpenStudy (jtug6):

yes. now you know what d is :)

OpenStudy (anonymous):

thanks a ton!

OpenStudy (jtug6):

mhm np. i generally ever get the chance to answer people and im not busy most of the time so jsut msg me directly if you ever need more help i dont care for the medals all too much :P